OpenStudy (anonymous):

A colony of bacteria grows exponentially. The colony begins with 8 bacteria, but 7 hours after the beginning of the experiment, it has grown to 720 bacteria. 1a. Give a formula for the number of bacteria as a function of time i got P(t) = 8(1.90)^(t/7), i divided 720/8 got 90 and did 90/100 for the percentage of growth value and thought it was 1.09, but its wrong

OpenStudy (anonymous):

please help

OpenStudy (anonymous):

Help please

myininaya (myininaya):

so at time t0=0, P(t0=0)=8 at time t7=7, P(t7=7)=720 one sec i brb

myininaya (myininaya):

are you still there?

myininaya (myininaya):

if you are... since we have an exponential thing going on we have something in this form P(t)=ab^t to find what a and b are we know two points on the curve (0,8) and (7,720)

myininaya (myininaya):

P(0)=ab^(0)=a1=a=8 since P(0)=8 so P(t)=8*b^t P(7)=8*b^(7)=720 since P(7)=720 we can solve this equation for b so we have 8b^7=720 b^7=720/8=90 b=90^(1/7) so P(t)=8*90^(t/7)